# This work has been supported by the Ministerio de Ciencia, Innovación y Universidades of the Spanish Government under HVD (PID2021-125051OB-I00) and SEGA-CV (TED2021-131643A-I00) projects, and the Comunidad Autónoma de Madrid under the Grant IND2020/TIC-17515

### v2

@@ Abstract
Data acquired in space operational conditions is scarce due to the costs and complexity of space operations. This poses a challenge to learning-based visual-based navigation algorithms employed in autonomous spacecraft navigation. Existing datasets, which largely depend on computer-simulated data, have partially filled this gap. However, the image generation tools they use are proprietary, which limits the evaluation of methods to unseen scenarios. Furthermore, these datasets provide limited ground-truth data, primarily focusing on the spacecraft’s translation and rotation relative to the camera. To address these limitations, we present SPIN (SPacecraft Imagery for Navigation), an open-source realistic spacecraft image generation tool for relative navigation between two spacecrafts. SPIN provides a wide variety of ground-truth data and allows researchers to employ custom 3D models of satellites, define specific camera-relative poses, and adjust various settings such as camera parameters and environmental illumination conditions. For the task of spacecraft pose estimation, we compare the results of training with a SPIN-generated dataset against existing synthetic datasets. We show a %50 average error reduction in common testbed data (that simulates realistic space conditions). Both the SPIN tool (and source code) and our enhanced version of the synthetic datasets will be publicly released upon paper acceptance on GitHub. https://github.com/vpulab/SPIN.
{IEEEkeywords}
Aerospace Navigation, Pose estimation, Simulation, Synthetic Data,
1 Introduction
@@
Visual-based navigation plays a crucial role in achieving autonomy, as it provides information on the elements of the environment and the relative position and orientation of the spacecraft. Current state-of-the-art methods for visual-based navigation employ learning techniques. However, the aforementioned lack of data prevents training robust learning-based navigation algorithms to support the required autonomy. To address this issue, researchers often turn to computer-based simulators that generate synthetic datasets, as detailed in Section 2. These datasets are essential for research in visual-based spacecraft navigation: they have enabled the development of algorithms capable of estimating spacecraft pose with centimetre-level accuracy [7] even in situations with large differences in scale [8].
@@
To bridge this gap, we present the SPacecraft Imagery for Navigation (SPIN) tool, the first open-source image generation tool designed specifically to create data for visual-based navigation between two spacecrafts. SPIN addresses the limitations of current state-of-the-art datasets, as illustrated in Figure 1, by enabling the generation of highly realistic and customizable spacecraft imagery in various poses or pose sequences. It also offers extensive ground truth data, encompassing depth, dense pose annotations, segmentation, and keypoints. SPIN allows to load any external spacecraft 3D model and comes pre-loaded with a model of the Tango spacecraft.
To validate SPIN, we show experiments results on the task of pose estimation, the common benchmark across all identified datasets (see Figure 1). To do so, we compare the results obtained from training with the SPIN-generated images against those obtained with two widely-used datasets: SPEED+ [7] and SHIRT [9] further introduced in Section  2. SPEED+ and SHIRT include both synthetic and testbed images –realistic imagery captured in laboratory conditions. We compare by training with a SPIN replication of the synthetic datasets, with the same poses but increased realism and ground-truth data, and then evaluating on the testbed images, yielding a 53% average reduction in error rate for spacecraft pose estimation tasks. We summarise our contributions as follows:
  • We provide the first open-source simulation tool designed to generate realistic datasets of spacecraft images along with depth, segmentation and dense pose ground-truth data.
  • Our tool allows to augment the ground-truth data of existing datasets, provided that the spacecraft’s 3D model is publicly available.
  • We provide an enhanced version of the existing SPEED+ [7] and SHIRT [9] datasets, with improved realism and additional depth, segmentation and dense pose labels.
2 Related Work
@@
We provide a detailed description of publicly available datasets based on monocular intensity images, summarising their features and limitations in Table 1. While our focus is on optical datasets, we acknowledge the existence of datasets derived from event sensors in the literature [10].
Table 1: Comparison of various established satellite pose estimation datasets with our simulation tool, SPIN. A checkmark ($\checkmark$) indicates the availability of the feature, while a dash (-) indicates its absence. Pose 6D refers to the pose encoded with the absolute position and rotation, whereas dense coordinates indicate the 3D position of each pixel of the spacecraft. The depth column represents the ground-truth metric depth, while the segmentation column represents the segmentation mask. The Images column indicates the availability of gray or colour images and their resolution. The Simulation column indicates the range of distances between spacecraft and camera covered by the dataset, and the availability of trajectories (video sequence).
@@
CubeSat-CDT [12] $\checkmark$ - - - RGB 1440x1080 [0.4m, 3.8m] $\checkmark$ $\checkmark$
@@
SPIN (Ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ RGB/Gray Configurable Configurable $\checkmark$ -
2.1 Datasets
@@
The SwissCube Dataset [8] constitutes a synthetic collection featuring 50,000 images of a 1U CubeSat model based on the SwissCube satellite. These images are organized into 500 trajectories, each comprising 100 frames, and are generated using the Mitsuba Renderer 2 framework. Importantly, the dataset encompasses a broad range of distances to the CubeSat and the camera, thereby introducing significant scale variability.
The CubeSat Cross-Domain Trajectory (CDT) dataset [12] comprises RGB images captured across multiple trajectories of a 1U CubeSat within three distinct domains. Specifically, the dataset includes two synthetic domains: the first, generated using Unity, contains 50 trajectories, and the second, created with Blender, encompasses 15 trajectories. Additionally, a testbed domain is provided, featuring 21 trajectories. The data from all three sources employ the same 1U CubeSat and share identical camera intrinsic parameters.
@@
The Next Generation Spacecraft Pose Estimation Dataset (SPEED+) [7], represented in the bottom row of Figure 2, consists of 60,000 synthetic images featuring the Tango spacecraft, accompanied by pose annotations. In addition, the dataset includes 9,531 annotated testbed images of a half-scale mock-up model. These test images are divided into two subsets: Sunlamp, which contains 2,791 images characterised by strong illumination and reflections against a dark background; and Lightbox, featuring 6,740 images with softer lighting conditions, elevated noise levels, and the presence of the Earth in the background.
The Satellite Hardware-In-the-loop Rendezvous Trajectories Dataset (SHIRT) [9], represented in the top row of Figure 2, features two distinct trajectories (video sequences), ROE1 and ROE2, capturing the poses of a Tango satellite from the perspective of a service spacecraft. Each sequence offers two sets of images: synthetic grayscale images and hardware-in-the-loop testbed images that are similar to the Lightbox subset in the SPEED+ dataset. Both the synthetic and testbed images capture the spacecraft in identical poses and are acquired using the same camera intrinsic parameters. This allows for the direct evaluation of the domain gap impact while holding all other variables constant.
We employ SPEED+ and SHIRT as benchmarks due to their relevance in the spacecraft pose estimation literature. SPEED+ was employed in the European Space Agency Spacecraft Pose Estimation Challenge 2021 [7]. SHIRT expands SPEED+ to contain sequences of images.
Figure 2: Representative images from the SHIRT [9] and SPEED+ [7] datasets. In the top row images from the SHIRT dataset: the image on the left is from the synthetic domain, while the one on the right is from the testbed domain. The bottom row contains images from the SPEED+ dataset, are arranged from left to right, showcasing the simulated and the testbed –Lightbox, and Sunlamp– settings.
2.2 Discussion
We argue that the current research in relative navigation between spacecrafts is constrained by existing datasets and the lack of simulation tools (Table 1). Firstly, datasets generally lack of diverse ground-truth labels, not providing depth, segmentation, or dense pose information. For instance, dense prediction techniques for spacecraft pose estimation have demonstrated their efficacy [15], while only one dataset provides the means to compute dense pose via its dense depth maps [11]. Additionally, self-supervised methods for estimating monocular depth and pose [16], are hard to assess due to a lack of datasets with trajectories (video sequences) containing ground-truth depth.
@@ 2.2 Discussion
3 SPIN: SPacecraft Imagery for Navigation
Figure 3 provides a schematic overview of our proposed tool, highlighting its essential features. The tool requires two primary inputs: a 3D model of a spacecraft (the Tango spacecraft is provided by default) and a set of predefined poses. Each pose defines the orientation and position of the spacecraft with respect to the camera so that the output is a set of images of the spacecraft in each defined pose with the associated ground-truth data.
Users have the flexibility to customise the rendered images acting on three main scene elements: 1) camera adjustments, specifically modifying intrinsic parameters, and sensor and lens imperfections: 2) altering the environment by tweaking illumination sources, background, and the rendering of shadows; and 3) activating or deactivating specific materials and material properties such as specularity.
The primary output of the tool are the spacecraft RGB or grayscale images, and the associated ground-truth data includes depth maps, dense pose maps, and segmentation masks. Additionally, a keypoint labelling tool to select keypoints and generate keypoint heatmaps is provided.
The tool has been developed in Unity Engine and will be distributed completely open-source.
3.1 Input
SPIN takes a 3D spacecraft model and a set of poses as inputs. We include a default 3D model of the Tango spacecraft, the one depicted in Figure 3. This model was created using the Fusion360 modelling software. There are multiple options for configuring spacecraft poses: manually setting the spacecraft’s location and rotation, importing camera-relative poses from a file, or generating random poses within the tool. In this last case, the tool employs the algorithm described in [20] to ensure uniform sampling of the rotation space.
3.2 Camera
The camera module governs image acquisition options and offers two configuration layers. The first layer focuses on lens distortion parameters and on the camera intrinsics, which determine parameters such as image size, focal length, and principal point. The second layer models specific challenges associated with space imaging, such as the absence of an atmosphere, low signal-to-noise ratio (SNR), expansive dynamic range, and highly reflective surfaces. These factors influence both the lens and the sensor. Within this layer, our tool provides options to simulate: a) sensor noise, b) lens glare and bloom coming from the sun and highly reflective surfaces, and c) different colour adjustment options to modify the overall colour temperature, intensity, and saturation of the final image.
3.3 Environment and Materials
The environment settings in our tool offer control over three main elements: illumination, background, and shadow rendering. For illumination, the tool includes three preset options: Spotlight, which simulates a spotlight in a fixed location and rotation with respect to the camera; Sunlight, offering directional lighting that replicates solar rays; and Ambient Light, producing diffuse illumination by combining multiple light sources around the target. All modes feature user-adjustable light intensity, the Spotlight and Sunlight options additionally allow to adjust rotation and position, and multiple Spotlight lights can be used and adjusted at the same time. Moreover, different illumination presets can also be enabled simultaneously.
As for the background, our tool not only supports a uniform backdrop but also includes a default selection of Earth images taken from space. These images can either be fixed or randomly positioned during each generation. Users have the flexibility to incorporate their own background images as well. Additionally, the tool provides the capability to toggle shadows and ambient occlusion on or off and allows for the customisation of various shadow properties.
Finally, the tool includes an option to enable or disable high-quality materials that offer enhanced textures and reflections. Given that spacecraft are often covered with reflective materials, this feature allows to choose non-Lambertian surfaces, which ensure consistent illumination across various spacecraft poses.
In Figure 4 we include some examples of images that can be generated using our simulator, with different settings to the ones used for the SPEED+ like images.
Figure 4: Some examples of different renderings generated with our tool. Top right image depicts the keypoint visualization tool, whereas the rest of the images present different combinations of background, illumination, and color corrections.
3.4 Outputs
The main output is the set of intensity images based on the configured camera, environment, and materials. Additionally, the tool can provide a metric depth map of the scene, a segmentation mask, and dense 3D coordinates of the spacecraft. While their coordinates can be derived from the depth map, we provide them separately to simplify the workflow and to ease research with prediction methods based on dense pose. In addition, we provide a way to generate keypoint heatmap for keypoint-based pose estimation tasks. This is detailed in Section 3.5.
3.5 Keypoint Tool
The tool includes a keypoint labelling module and a keypoint heatmap generation module. Keypoint-based methods currently represent the state-of-the-art in spacecraft pose estimation [6]. These methods rely on estimating the 2D projection of predefined 3D keypoints (as heatmaps) on the spacecraft and solving for 2D-3D correspondences to determine the pose. Despite their effectiveness, few datasets provide a set of keypoints. The typical process involves selecting a set of 2D points, triangulating them, and further optimising the triangulation using convex solvers [21]. As we are aware that this is a time-consuming task, to facilitate research we provide a labelling mechanism for keypoints and a heatmap generation module to create the heatmap representation of the 2D keypoints.
Table 2: Summary of available parameters and options in the simulator. The two right-most columns indicate which parameters have been modified for increased realism and which ones have been set to preserve a consistent input domain w.r.t the original synthetic split of SPEED+ and SHIRT datasets.
Module Parameter Description Options Modification w.r.t SPEED+ SHIRT Justification
Camera Intrinsics + Lens distortion Camera sensor and lens configuration Sensor Size, FOV, ISO, Aperture… No Preserve consistent input domain
Color Adjustment Enables different color adjustment options over the generated image Saturation, Contrast, Hue Shift… No Preserve consistent input domain
Noise Simulates different types of sensor noise Noise type, intensity, response. No Preserve consistent input domain
Glare/Bloom Simulates camera glare effects. Threshold, Intensity, Scatter, Tint… Yes Enhance scene realism to align with testbed conditions
Environment Illumination Multiple illumination options. Ambiental Illumination, Directional, Spotlight, Light intensity, Light temperature… Yes Enhance scene realism to align with testbed conditions
Background Allows the user to set up background images. Background image, position, rotation, random positions. No Maintain a consistent Earth background across respective poses for uniform impact
Shadows Enables shadows on the simulator. Shadow quality, Shadow distance, Ambiental Occlusion… No Retain original shadow rendering
Material High-quality materials Enable or disable high-quality materials for the satellite. Material quality, Specularity on/off. Yes Enhance scene realism to align with testbed conditions
4 Experimental Validation
@@
We propose to validate SPIN by employing its output to train a model for the task of spacecraft pose estimation, the common benchmark across the existing datasets. More in detail, we use SPIN to replicate the synthetic images of SPEED+ [7] and SHIRT [9] for our validation (examples provided in Figure 2). SPEED+ shows the particularity of having a synthetic and two testbed domains which, in addition to the testbed domain of SHIRT, provide a suitable framework for evaluating the quality of SPIN and its impact on sim-to-real transfer. The experiments are organised as follows, we first describe the settings for the experiments on the pose estimation. Next, we detail the configurations used in SPIN for image generation. We conclude by presenting the quantitative results, including an ablation study.
4.1 Pose Estimation Experimental Settings

### v3

# SPIN: An Open Simulator of Realistic Spacecraft Navigation Imagery

@@ Abstract
The scarcity of data acquired under actual space operational conditions poses a significant challenge for developing learning-based visual navigation algorithms crucial for autonomous spacecraft navigation. This data shortage is primarily due to the prohibitive costs and inherent complexities of space operations. While existing datasets, predominantly relying on computer-simulated data, have partially addressed this gap, they present notable limitations. Firstly, these datasets often utilize proprietary image generation tools, restricting the evaluation of navigation methods in novel, unseen scenarios. Secondly, they provide limited ground-truth data, typically focusing solely on the spacecraft’s translation and rotation relative to the camera. To address these limitations, we present SPIN (SPacecraft Imagery for Navigation), an open-source spacecraft image generation tool designed to support a wide range of visual navigation scenarios in space, with a particular focus on relative navigation tasks. SPIN provides multiple modalities of ground-truth data and allows researchers to employ custom 3D models of satellites, define specific camera-relative poses, and adjust settings such as camera parameters or environmental illumination conditions. We also propose a method for exploiting our tool as a data augmentation module. We validate our tool on the spacecraft pose estimation task by training with a SPIN-generated replica of SPEED+[1], reaching a 47% average error reduction on SPEED+ testbed data (that simulates realistic space conditions), further reducing it to a 60% error reduction when using SPIN as a data augmentation method. Both the SPIN tool (and source code) and our SPIN-generated version of SPEED+ will be publicly released upon paper acceptance on GitHub. https://github.com/vpulab/SPIN.
I Introduction
@@
Visual-based navigation plays a crucial role in achieving autonomy, as it provides information on the elements of the environment and the relative position and orientation of the spacecraft. Current state-of-the-art methods for visual-based navigation employ learning techniques. However, the aforementioned lack of data prevents training robust learning-based navigation algorithms to support the required autonomy. To address this issue, researchers often turn to computer-based simulators that generate synthetic datasets, as detailed in Section II. These datasets are essential for research in visual-based spacecraft navigation: they have enabled the development of algorithms capable of estimating spacecraft pose with centimetre-level accuracy [1] even in situations with large differences in scale [8].
@@
To bridge this gap, we present the SPacecraft Imagery for Navigation (SPIN) tool, the first open-source image generation tool designed specifically to create data for visual-based navigation between two spacecrafts. SPIN addresses the limitations of current state-of-the-art datasets, as illustrated in Fig. 1, by enabling the generation of highly realistic and customizable spacecraft imagery in various poses or pose sequences. It also offers extensive ground truth data, encompassing depth, dense pose annotations, segmentation, and keypoints. SPIN allows to load any external spacecraft 3D model and comes pre-loaded with a model of the Tango spacecraft.
We validate SPIN performing experiments on the task of pose estimation, the common benchmark across all identified datasets (see Fig. 1). We compare the results obtained from training with the SPIN-generated images against those obtained the widely used dataset SPEED+ [1], and evaluate on the testbed images –realistic imagery captured in laboratory conditions– from SPEED+ and SHIRT [9] further introduced in Section  II. Training with our SPIN-generated replica from the SPEED+ dataset, yields a 47% average reduction in error rate for spacecraft pose estimation tasks, which is further improved to an average of 60% error rate reduction when performing data augmentation using our simulation tool.
We summarise our contributions as follows:
  • We provide the first open-source simulation tool designed streamline the generation of datasets for spacecraft imagery along with depth, segmentation and dense pose ground-truth data.
  • We propose a novel method to leverage our simulator, in-scene data augmentation, by augmenting the actual environment and not just the 2D image.
  • We share an enhanced version of the existing SPEED+ [1] dataset with additional depth, segmentation and dense pose labels, generated using our simulation tool.
II Related Work
@@
We provide a detailed description of publicly available datasets based on monocular intensity images, summarising their features and limitations in Table I. While our focus is on optical datasets, we acknowledge the existence of datasets derived from event sensors in the literature [10].
TABLE I: Comparison of various established satellite pose estimation datasets with our simulation tool, SPIN. A checkmark ($\checkmark$) indicates the availability of the feature, while a dash (-) indicates its absence.
@@
CubeSat [12] $\checkmark$ - - - RGB 1440x1080 [0.4m, 3.8m] $\checkmark$ $\checkmark$
@@
SPIN (Ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ RGB/Gray Custom Custom $\checkmark$ -
II-A Datasets
@@
The SwissCube Dataset [8] constitutes a synthetic collection featuring 50,000 images of a 1U CubeSat model based on the SwissCube satellite. These images are organized into 500 trajectories, each comprising 100 frames, and are generated using the Mitsuba Renderer 2 framework.
The CubeSat Cross-Domain Trajectory (CDT) dataset [12] comprises RGB images captured across multiple trajectories of a 1U CubeSat within three distinct domains. Specifically, the dataset includes two synthetic domains: the first, generated using Unity, contains 50 trajectories, and the second, created with Blender, encompasses 15 trajectories. Additionally, a testbed domain is provided, featuring 21 trajectories.
@@
The Next Generation Spacecraft Pose Estimation Dataset (SPEED+) [1], represented in the bottom row of Fig. 2, consists of 60,000 synthetic images featuring the Tango spacecraft, accompanied by pose annotations. In addition, the dataset includes 9,531 annotated testbed images of a half-scale mock-up model. These test images are divided into two subsets: Sunlamp, which contains 2,791 images characterised by strong illumination and reflections against a dark background; and Lightbox, featuring 6,740 images with softer lighting conditions, elevated noise levels, and the presence of the Earth in the background.
The Satellite Hardware-In-the-loop Rendezvous Trajectories Dataset (SHIRT) [9], represented in the top row of Fig. 2, features two distinct trajectories (video sequences), ROE1 and ROE2, capturing the poses of a Tango satellite from the perspective of a service spacecraft. Each sequence offers two sets of images: synthetic grayscale images and hardware-in-the-loop testbed images that are similar to the Lightbox subset in the SPEED+ dataset.
In this work we SPEED+ and SHIRT as benchmarks due to their relevance in the spacecraft pose estimation literature. SPEED+ was employed in the European Space Agency Spacecraft Pose Estimation Challenge 2021 [1]. SHIRT expands SPEED+ to contain sequences of images.
II-B Discussion
We argue that the current research in relative navigation between spacecrafts is constrained by existing datasets and the lack of simulation tools (Table I). Firstly, datasets generally lack of diverse ground-truth labels, not providing depth, segmentation, or dense pose information. For instance, dense prediction techniques for spacecraft pose estimation have demonstrated their efficacy [15], while only one dataset provides the means to compute dense pose via its dense depth maps [11]. Additionally, self-supervised methods for estimating monocular depth and pose [16], are hard to assess due to a lack of datasets with trajectories (video sequences) containing ground-truth depth.
Figure 2: Representative images from the SHIRT [9] and SPEED+ [1] datasets. Top row are from SHIRT dataset: Left image - synthetic domain; Right image - testbed domain. Bottom row are from SPEED+ dataset: Left to right - simulated, testbed Lightbox, testbed Sunlamp settings.
@@ II-B Discussion
III SPIN: SPacecraft Imagery for Navigation
Fig. 3 provides a schematic overview of our proposed tool, showing its essential features, an providing an outline on different adjustable parameters from our simulator. Additionally, along the simulator we include tools to aid in the task of keypoint labeling, and also keypoint heatmap generation.
III-A Simulation Framework
Our tool has been built relying on Unity Engine, which has also been previously used in related literature (CubeSat[12] and SPARK [11]). Unity Engine includes multiple options for Cameras, Lightning, Materials, as well as different environment settings, and post-processing effects.
Cameras in Unity Engine are highly configurable, and include a vast array of parameters to imitate a physical camera. From different sensor parameters such as ISO, Sensor size or pixel size, to Lens parameters such as aperture, focal length and distance.
Unity provides different types of object-source lightning, where you can adjust some parameters such as light intensity, temperature, and whether or not the light source must cast shadows. Also, the source of light can be set up with lightning types: Directional light simulates sunlight, as if the source of light was at an infinite distance, and allows the user to set up the angular diameter for such light. Spotlight simulates a focused light aimed at a point, and user can set up the shape and reach of the spotlight, Area light simulates a soft box light , were light comes uniformly from an area and not a point in space, and the user can set up the area shape and size, and finally Point light generates light in all directions from a given point, to a maximum radius set up by the user.
Materials in the Unity editor are customizable and allow the user to include different textures for albedo (as RGB), specular reflections, masks, and parameters such as self-emission, and textures such as normal maps and displacement maps that can be used to simulate fine detail without the computational overhead of additional polygons in the satellite model. Unity provides different shaders designed using Phisically-Based-Rendering, which we use in our simulator due to ease of use and perceived quality.
The main advantage of tools such as Unity Engine is how they enable setting up virtually any kind of environment, combining the elements previously mentioned along with 3D models, to simulate different sceneries. For this purpose, although a user may virtually be able to modify everything, it does require coding and more in-depth knowledge of specific Unity Engine libraries and elements.
In our framework, we provide an easy, straightforward way to modify these parameters and setup environments adjusted to spacecraft imagery generation that is also easy to use and does not require specific Unity Engine knowledge or experience, so researchers that may not be familiar with 3D engines can still generate their own data adjusted to their tasks.
III-B Input and Output Data
Users can set up different environments in the simulator by setting up 3D model and poses relative to camera. By default, we include a Tango spacecraft model we built using the Speed [14] dataset as reference, but users can also include their own 3D models and set up its materials and dimensions.
To set up camera locations, user can provide a list of relative camera poses, following the format from Speed+ [1], describing the relative position using a quaternion $q={x,y,z,w}$ and a translation vector $v={x,y,z}$.
Users can generate different modalities of data: RGB or Grayscale images, depth, dense pose or LiDAR, and ground-truth for tasks such as object detection (both 2D and 3D bounding boxes), semantic segmentation or occlusion masks. We use the Perception Unity Package [20] to generate these multimodal data and ground-truths.
III-C Additional tools
Additionally, we include a Keypoint tool to allows users to easily extract an accuracte keypoint list for different 3D models. Related to this tool we provide another tool to generate keypoint heatmaps (methods that use keypoint heatmaps represent the state-of-the-art in spacecraft pose estimation [7]) for a given dataset using these user-generated keypoints.
IV Simulator Validation
@@
We propose to validate SPIN by employing its output to train a model for the task of spacecraft pose estimation, the common benchmark across the existing datasets. More in detail, we use SPIN to replicate the synthetic images of SPEED+ [1] and SHIRT [9] for our validation (examples provided in Fig. 2). SPEED+ shows the particularity of having a synthetic and two testbed domains which, in addition to the testbed domain of SHIRT, provide a suitable framework for evaluating the quality of SPIN and its impact on sim-to-real transfer. The experiments are organised as follows, we first describe the settings for the experiments on the pose estimation. Next, we detail the configurations used in SPIN for image generation. We conclude our validation by presenting the quantitative results, including an ablation study.
IV-A Pose Estimation Experimental Settings